OpenStudy (anonymous):

Find the equation for the line that passes through point (3,4) and is parallel to the equation 3/2x-3y=27/2

OpenStudy (amistre64):

parallel means you use the same equation just plug in your x and y values to get a new sum

OpenStudy (amistre64):

we can also start by getting rid of the fractions; multiply it all by 2

OpenStudy (amistre64):

3x -6y = ..... ; now use your x and y

OpenStudy (anonymous):

im still confussed...

OpenStudy (amistre64):

thats becasue you havent learned how math works yet...

OpenStudy (amistre64):

if you tell me what your confused about, i might be able to help

OpenStudy (anonymous):

nope i slept through my classes in high school and now its coming back to haunt me

OpenStudy (amistre64):

do you recall the 5/5 = 1?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then thats all i did to clear the fractions: if theres a 2 on the bottom, what needs to be on top to clear it? a 2 right? 2/2 = 1

OpenStudy (amistre64):

whatever you do to one thing you do to them all: 3(2)/2x -3(2)y = 27(2)/2 3x -6y = 27 looks better right?

OpenStudy (anonymous):

i can do english or other subjects no problem but this stuff makes me feel so stupid

OpenStudy (amistre64):

do you understand how i cleared the fractions?

OpenStudy (anonymous):

yes that looks better

OpenStudy (anonymous):

kinda

OpenStudy (amistre64):

now, to get a parallel line; we use the same equation but that 27 is garbage to us; throw it out and get a new number there

OpenStudy (amistre64):

this point they give you: (3,4) MEANS (x=3,y=4), so wherever you see an x plug in a 3; wherever you see a y plug in a 4

OpenStudy (anonymous):

k...3x-6y so 3(3)-(4) ?

OpenStudy (amistre64):

3x -6y = ____ 3(3) - 6(4) = ____ 9 - 24 = -15 ; thats our new number, -15 3x -6y = -15

OpenStudy (anonymous):

nooooooooo ok thats wat i was trying to type

OpenStudy (amistre64):

:) typing and mathing are not compatible lol

OpenStudy (anonymous):

ok im starting to kinda get it

OpenStudy (anonymous):

im figuring that out

OpenStudy (anonymous):

ok so now that we have this new equation how do i put it into slope intercept form

OpenStudy (amistre64):

slope intercept means move it all over to one side except for a y

OpenStudy (anonymous):

right

OpenStudy (amistre64):

3x -6y = -15 ; -3x from both sides -6y = -3x -15 ; multiply it all by -1 6y = 3x +15 ; divide both sides by 6 y = (3/6)x + 15/6 ; reduce fractions y = (1/2)x + 5/2

OpenStudy (anonymous):

where did u get 1 to multiply it by

OpenStudy (anonymous):

-1 i mean

OpenStudy (amistre64):

I could have used 128 and it would have changed nothing of the value.... but that seemed excessive lol

OpenStudy (amistre64):

tell me does: 3 = 3?

OpenStudy (anonymous):

ok...so it doesnt matter wat u multiply it by yes

OpenStudy (amistre64):

since 3=3 can we multiply at all by 4 and get a true result? 3=3 [4] 12 = 12 ??

OpenStudy (anonymous):

um yes

OpenStudy (amistre64):

then it really doesnt matter "where" I got it from, cause the results are the same regardless right?

OpenStudy (anonymous):

yes i see now

OpenStudy (amistre64):

I choose it because its useful to me , thats all.

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

I wanted to clear out the negatives; and used 1 to keep the number looking the same

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

-3x = -14 ; if we multiply it all by -1, we clear out the negative signs 3x = 14 ; now we have a prettier equation to work with and it still has the same value

OpenStudy (anonymous):

when i look at this its like another language i cant understand.

OpenStudy (amistre64):

:) thats cause math had to invent new symbols to express itself

OpenStudy (amistre64):

the only place you see 1+1 is on paper lol

OpenStudy (anonymous):

im glad that there are people who can understand it

OpenStudy (amistre64):

me too ;)

OpenStudy (anonymous):

im so mathematically challenged its not even funny

OpenStudy (anonymous):

the fact that this class is partly online doesnt help either

OpenStudy (amistre64):

im sure you mean doing it on paper; math is something you do all the time, its just that when we try to define it so that we can discuss it, people get wierded out

OpenStudy (anonymous):

its called hawkins courseware. it doesnt explain it very clearly

OpenStudy (amistre64):

i tried an online course once, had to drop it cause my car broke and couldnt afford books afterwards, but id much rather sit in a class

OpenStudy (anonymous):

i can do english or history online but im not so sure i can do this algebra

OpenStudy (amistre64):

:) algebra is fine

OpenStudy (anonymous):

i hope one day i will think that way...lol

OpenStudy (anonymous):

thanks for your help

OpenStudy (amistre64):

youre welcome :)